How do i solve a system of equations by graphing such as
7x-7y=35
7x+7y=-35

Answer by richwmiller(17219)

They intersect at x=0 y=-5
plug in y=0 in the first and get x=5
plug in y=0 in the second and x=-5
plug in x=0 and get y=-5 for both
if x=1 7-7y=35 -7y=28 y=-4
if x=1 7+7y=-35 7y=-42 y=-6
plot those points and draw the lines and see that they cross at x=0 y=-5